Write down gauses theorem and stablish coulombs inverse square law by using this theorem



1
Expert's answer
2021-03-26T11:39:00-0400

Write down Gauss's theorem as he wrote it:


"\int\bigg(\frac{dP}{dx}+\frac{dQ}{dy}+\frac{dR}{dz}\bigg)d\omega=\int(P\cos\alpha+Q\cos\beta+R\cos\gamma)ds"


However, another formula is famous in the field of electromagnetism:


"\oint\vec{E}\vec{dS}=\frac{Q_\text{net}}{\epsilon_0}."

Obtain inverse square law. For the electric field from a point charge, we have


"\oint\vec{E}\vec{dS}=\int EdS\cos0°,\\\space\\
\oint\vec{E}\vec{dS}=E\int dS=E(4\pi R^2),\\\space\\"

with Gauss's theorem we have


"\oint\vec{E}\vec{dS}=\frac{Q_\text{net}}{\epsilon_0},\\\space\\
E(4\pi R^2)=\frac{Q_\text{net}}{\epsilon_0},\\\space\\
E=\frac{1}{4\pi\epsilon_0}\frac{Q}{R^2}."